# A Large Scale Analysis of Gender Biases in Text-to-Image Generative Models

### v1

@@ Abstract
With the increasing use of image generation technology, understanding its social biases, including gender bias, is essential. This paper presents the first large-scale study on gender bias in text-to-image (T2I) models, focusing on everyday situations. While previous research has examined biases in occupations, we extend this analysis to gender associations in daily activities, objects, and contexts. We create a dataset of 3,217 gender-neutral prompts and generate 200 images per prompt from five leading T2I models. We automatically detect the perceived gender of people in the generated images and filter out images with no person or multiple people of different genders, leaving 2,293,295 images. To enable a broad analysis of gender bias in T2I models, we group prompts into semantically similar concepts and calculate the proportion of male- and female-gendered images for each prompt. Our analysis shows that T2I models reinforce traditional gender roles, reflect common gender stereotypes in household roles, and underrepresent women in financial-related activities. Women are predominantly portrayed in care- and human-centered scenarios, and men in technical or physical labor scenarios. Code and data will be released.
Figure 1: Visualization of our image generation setup to analyze gender bias in T2I models with respect to everyday scenarios.
@@ 1 Introduction
Rapid advances in image generation technology make it easier than ever to automatically generate large amounts of synthetic images. State-of-the-art text-to-image (T2I) models [52, 83] can generate high-quality images from arbitrary text instructions. Their capabilities are further enhanced through editing [10, 49, 101] and personalization [75, 34, 50, 9] techniques. Synthetic images are not only used in everyday applications such as advertisements [53] and presentation slides [67] but are also increasingly used as training data for other foundation models [86, 87, 30, 50].
@@
As the availability and proliferation of synthetic images increase, so does their power to influence society and amplify any harms originating from the underlying models [15]. In their seminal work, [12] discovered intersectional gender and racial biases in image recognition systems. Within the research community, the list of known biases has only grown: [5, 38, 7, 81, 84] identified social biases in CLIP [69], such as associating men with words related to criminal activities. [40, 45, 46] found social biases in automatic image captioning. [100, 96, 36, 32, 74] uncovered various social biases in multimodal large language models (MLLMs). These examples demonstrate that social bias pervades all aspects of modern generative AI systems.
@@
Research on social bias in T2I models has led to a large body of work covering all computational aspects of social bias, including bias analysis [21, 8, 56], open bias detection [24, 20, 23], and model debiasing [99, 6, 28, 33]. However, most research in this area has focused on gender-occupation bias [89]. While this is an important issue, other aspects of daily life, such as everyday activities and stereotypical contexts, also contribute to perpetuating or amplifying harmful social biases and require careful analysis. Furthermore, many studies on social bias in T2I models use a limited set of prompts and only a few images per prompt, reducing their representativeness.
@@
This work presents the first large-scale study on gender bias in T2I models concerning everyday contexts. We address gaps in the literature by analyzing everyday activities complemented by gender-object and gender-context associations. Our main contributions are: (1) We compile 3,217 gender-neutral prompts over four categories to probe T2I models for gender bias as illustrated in Fig. 1; (2) We generate 200 images from five state-of-the-art T2I models for each prompt and filter unsuitable images, leaving 2,293,295 images for analysis; (3) We design a carefully structured experimental setup to analyze gender bias in large image datasets as detailed in Fig. 2; (4) We systematically examine the observed gender biases and relate them to known human gender biases. Through this work, we aim to take an important step toward addressing gender stereotypes in T2I models, thereby helping to understand perpetuation of gender inequality in generative AI technology [85, 2].
@@ 2 Related Work
Prior work on evaluating social bias in text-to-image models can be categorized into bias analysis and open bias detection. Notable works on open bias detection include TIBET [20], OpenBias [24], and OASIS [23]. In this work, we conduct a careful, large-scale study of gender bias in state-of-the-art diffusion models. This paper focuses specifically on analyzing gender bias in text-to-image models by exploring gendered outputs in gender-neutral prompts.
Work analyzing social bias in text-to-image (T2I) generation has traditionally focused on a few categories [89], most notably perceived gender and race, which are protected against discrimination under U.S. federal law. In their seminal work, [8] find that T2I models amplify bias across many categories. [19] examine gender and racial bias in occupations and report that generated images are more gender-imbalanced than actual U.S. labor statistics. However, [80] provides evidence that bias amplification may stem from specific prompt design choices. [56] propose a method to analyze gender and racial bias without explicitly labeling images by gender or race, which can be controversial due to assigning gender and race to images. They find distributional differences in images generated for occupation-related prompts. Similarly, [21] observe distributional differences when generating images of occupations using gendered and gender-neutral prompts. Other studies explore different aspects of bias. [102, 93, 59] investigate gender bias in person-object co-occurrence. [88] find that non-binary identities are poorly represented in T2I models. Likewise, [35] show that T2I models inadequately represent national identities and tend to produce overly sexualized images of women, particularly when prompted for regions in the Global South. Expanding this perspective, [48] quantifies national stereotyping in generated images.
Previous work has established that T2I models exhibit gender bias, among other issues. However, as also pointed out by [89], scenarios beyond occupation are underrepresented in current research. Furthermore, most studies are based on small-scale analyses, typically using fewer than 200 prompts and no more than 20 images per prompt. Benchmarks that aim for a larger scale, such as [57], are still dominated by occupation-related prompts. Thus, this work provides the means for large-scale analyses of gender bias beyond occupations. Additionally, [58] find that many images generated from commonly used prompt templates do not depict people with distinguishable gender characteristics, so we ensure the validity of our analysis by filtering images without people or showing both men and women.
@@ 3.1 Prompt Collection, Processing and Clustering
We collect prompts in four categories: (1) Activities, (2) Contexts, (3) Objects, and (4) Occupations. To study gender bias in everyday activities, we rely on the curated set from [92], i.e. Activities. The activities in [92] were gathered through Amazon Mechanical Turk from U.S. based workers, who were asked to provide short phrases describing recent activities. Including these 1405 activities provides insight into how stereotypical gender roles in everyday life are reflected in T2I models. Analyzing gender associations with contexts and objects further extends this analysis by considering everyday situations beyond specific activities. To study gender bias in relation to people and places, i.e. Contexts, we collect a set of 737 scene classes from the SUN Database [94, 95]. We include all classes but do not consider fine-grained distinctions, e.g. “inside” the church and “outside” the church leads to the context “church”.
We collect 500 common physical objects from WordNet [31], i.e. Objects. To select these 500 objects, we filter all noun hyponyms of the object.n.01 WordNet synset using a list of the most common English words. Additionally, we manually remove a small number of unsuitable synsets, e.g. synsets that refer to people or body parts, or places that are already covered in the contexts prompt group. We retain the top 1000 most frequent lemmas and re-rank them based on their concreteness following [11]. Finally, we select the top 500 most concrete lemmas from the top 1000 most frequent WordNet lemmas. We also include occupations to align with prior work and to examine occupation-related gender bias in T2I models. Our occupation list is comparatively large, as we include all 575 occupations listed by the U.S. BLS [64] rather than a subset, i.e. Occupations.
Using an LLM, specifically Yi-1.5-34B [98] (see Sec. A.2 for a comparison to other LLMs), we process all collected activities, contexts, objects, and occupations to generate syntactically coherent prompts. Since we do not specify gender in our prompts, each prompt begins with “a person”. We then apply a different template for each prompt group as shown in Tab. 1. Detailed prompts for filling in the templates can be found in Sec. A.1. Additionally, we simplify occupation descriptions, which are often overly detailed in [64].
Activity a person who is {{verb}}-ing {{activity}}”
Context a person {{in/on/at…}} the {{context}}”
Objects a person and {{a object}}”
Occupations a person working as {{occupation}}”
@@
We generate 5 variations per prompt by replacing the prefix “a person” with “an individual”, “someone”, “a friend”, and “a colleague”, which do not include any gender information. Prompt variations increase diversity and are essential for valid analysis of social bias in large models [79, 78, 42]. Sec. B.3 shows that variations do not lead to a significant skew towards male- or female-gendered images.
@@
For a concise presentation of our findings from the 3217 prompts, we cluster prompts in all four prompt groups into semantically coherent concepts to facilitate analysis. We apply the following variation of BERTopic [37] to cluster prompts. First, we embed all prompts using a sentence embedding model [70]. Then, we reduce the embedding dimensions to 16 using UMAP [62]. Using HDBSCAN clustering [61] with cosine distance, we determine the concept clusters. However, HDBSCAN has the option to not assign a prompt to any cluster, so we add unclustered prompts to the cluster with nearest centroid. The detailed settings are in Sec. A.3. We arrive at 165 activity clusters, 91 context clusters, 62 object clusters, and 76 occupation clusters.
After clustering prompts, we summarize all clusters by an LLM (see Sec. A.3). Summarizing a list of prompts requires more in-depth understanding and reasoning than prompt processing, so we use Llama-3-3-70B-Instruct based on manual inspection. A comparison to other LLMs and our exact prompt template is in Sec. A.3. For the purposes of our analysis, we further merge clusters with the same summary (e.g. different variants of shopping-related activities), as they would be indistinguishable for the reader. However, we will make the fine-grained clusters available alongside our code.
@@ 3.2 Image Generation and Gender Identification
We generate images using 5 models, which represent the state-of-the-art among open models at the time of writing: (1) Flux [52], (2) Flux-Schnell, (3) Stable Diffusion 3.5 Large [83], (4) Stable Diffusion 3.5 Medium, and (5) Stable Diffusion 3 Medium [29]. These are latent diffusion models [72] based on Diffusion Transformers [66]. Other recent strong models, such as Lumina 2.0 [55] and Janus-Pro-7B [16], were concurrently released to this study and unavailable when conducting experiments. For each combination of model and prompt, we generate 40 images per prompt variation. This results in 5 models $\times$ 5 prompt variations $\times$ 3217 prompts $\times$ 40 images = 3,217,000 images.
To analyze gender bias in generated images, we identify the perceived gender of the people shown in the images using a two-step process. First, we detect all people in the generated images using the object detection model YOLOv10 [90] and obtain bounding boxes for the detected individuals. Next, we crop each detected person’s bounding box and pass it to an MLLM, InternVL2-8B [17, 18], along with a prompt asking the model to identify the person’s gender as “female”, “male”, or “unclear/cannot tell”. We focus on binary perceived gender, specifically men and women, for several reasons. First, current T2I models do not generate distinct features that clearly indicate non-binary gender. Second, current MLLMs rarely consider non-binary gender as an option, as shown in Sec. C.2. While not addressed in this paper, the fact that models output gender as binary is a separate issue that warrants further discussion.
Using bounding boxes instead of the full image helps mitigate bias from the person’s context, i.e. predicting the gender based on the background and not the person’s features, as MLLMs also exhibit gender bias [36]. It also prevents confusion when multiple people of potentially different genders appear in the image. In Sec. C.1, we provide the detailed prompt used for gender identification and verify, using human-labeled data, that the MLLM used in this study can identify perceived gender with near-perfect accuracy. It is important to note that assigning a person’s gender can be problematic, because it cannot necessarily be perceived from an image, and gender is a spectrum. Therefore, good practice with images of real people is to have people self-identify their gender. However, T2I models create images that are not real, so assigning perceived gender to the images is more acceptable as there is no risk of misidentifying a real person.
@@
Table 2: Prompt Groups with remaining prompts after filtering, and the exact number of remaining images in brackets.
@@ 3.3 Image and Prompt Filtering
We exclude images and prompts that are not suitable for analyzing gender bias, i.e. if (a) There is no person in the image; (b) There is no person in the image whose gender can be clearly identified (see Sec. C.3 for details); (c) There are multiple people in the image and there is at least one man and one woman. If there are multiple people, we keep the image if people with all the same gender are shown or where the gender of other people is labeled “unclear/cannot tell.” (for example people in the background).
Additionally, we exclude entire prompts for a model if fewer than 100 out of 200 images remain after filtering. Since we analyze gender bias at the prompt level, we can only consider prompts where we can reliably estimate the ratio of male and female people in the images. The number of remaining prompts for each model is in Tab. 2.
Figure 4: (Top) Top 10 most female-dominated (top row) and top 10 most male-dominated (bottom row) activity clusters. Bars indicate the ratio of female-gendered images generated from 5 T2I models averaged over prompts in each cluster, and the error line indicates the standard deviation across prompts. (Bottom) Top 5 most female-dominated (left) and top 5 most male-dominated (right) context clusters.

### v2

@@ Abstract
With the increasing use of image generation technology, understanding its social biases, including gender bias, is essential. This paper presents a large-scale study on gender bias in text-to-image (T2I) models, focusing on everyday situations. While previous research has examined biases in occupations, we extend this analysis to gender associations in daily activities, objects, and contexts. We create a dataset of 3,217 gender-neutral prompts and generate 200 images over 5 prompt variations per prompt from five leading T2I models. We automatically detect the perceived gender of people in the generated images and filter out images with no person or multiple people of different genders, leaving 2,293,295 images. To enable a broad analysis of gender bias in T2I models, we group prompts into semantically similar concepts and calculate the proportion of male- and female-gendered images for each prompt. Our analysis shows that T2I models reinforce traditional gender roles and reflect common gender stereotypes in household roles. Women are predominantly portrayed in care and human-centered scenarios, and men in technical or physical labor scenarios.
@@ 1 Introduction
Rapid advances in image generation technology make it easier than ever to automatically generate large amounts of synthetic images. State-of-the-art text-to-image (T2I) models (Black Forest Labs, 2024; Stability AI, 2024) can generate high-quality images from arbitrary text instructions. Their capabilities are further enhanced through editing (Brooks et al., 2023; Kawar et al., 2023; Zhang et al., 2024b) and personalization (Ruiz et al., 2023; Gal et al., 2023; Kim et al., 2024; Bini et al., 2024) techniques. Synthetic images are not only used in everyday applications such as advertisements (Lin et al., 2023) and presentation slides (Peng et al., 2024) but are also increasingly used as training data for other foundation models (Tian et al., 2024a; b; Fan et al., 2024; Kim et al., 2024). Here, Chen et al. (2024a) have observed age and skin tone bias amplification at increased levels of synthetic images in the pretraining data.
@@
As the availability and proliferation of synthetic images increase, so does their power to influence society and amplify any harms originating from the underlying models (Chan et al., 2023). In their seminal work, Buolamwini & Gebru (2018) discovered intersectional gender and racial biases in image recognition systems. Within the research community, the list of known biases has only grown: Agarwal et al. (2021); Hall et al. (2024); Berg et al. (2022); Seth et al. (2023); Tanjim et al. (2024) identified social biases in CLIP (Radford et al., 2021), such as associating men with words related to criminal activities. Hendricks et al. (2018); Hirota et al. (2022; 2023) found social biases in automatic image captioning. Zhang et al. (2024a); Xiao et al. (2024); Girrbach et al. (2025); Fraser & Kiritchenko (2024); Ruggeri et al. (2023) uncovered various social biases in multimodal large language models (MLLMs). These examples demonstrate that social bias pervades all aspects of modern generative AI systems.
@@
Research on social bias in T2I models has led to a large body of work covering all computational aspects of social bias, including bias analysis (Cho et al., 2023; Bianchi et al., 2023; Luccioni et al., 2024), open bias detection (D’Incà et al., 2024; Chinchure et al., 2024; Dehdashtian et al., 2025), and model debiasing (Zhang et al., 2023; Bansal et al., 2022; Esposito et al., 2023; Friedrich et al., 2024). However, most research in this area has focused on gender-occupation bias (Wan et al., 2024). While this is an important issue, other aspects of daily life, such as everyday activities and stereotypical contexts, also contribute to perpetuating or amplifying harmful social biases and require careful analysis. Furthermore, many studies on social bias in T2I models use a limited set of prompts and only a few images per prompt, reducing their representativeness.
@@
We present a large-scale, in-depth study on gender bias in T2I models concerning everyday scenrios and address gaps in the literature by analyzing everyday activities complemented by gender-object and gender-context associations. Our main contributions are: (1) We compile 3,217 gender-neutral prompts over four categories to probe T2I models for gender bias and generate 200 images from five state-of-the-art T2I models for each prompt and filter unsuitable images, leaving 2,293,295 images for analysis; (2) We design a carefully structured experimental setup to analyze gender bias in large image datasets and systematically examine the observed gender biases and relate them to known human gender biases; (3) We analyze bias amplification in activities compared to LAION-400m and confirm bias amplification in occupations wrt. U.S. labor statistics. Our work takes an important step toward addressing gender stereotypes in T2I models, helping understand and document perpetuation of gender inequality in AI technology (Tannenbaum et al., 2019; Abebe et al., 2020).
@@ 2 Related Work
Work analyzing social bias in T2I generation has largely targeted a few categories (Wan et al., 2024), especially perceived gender and race, and mostly used occupation prompts. In their seminal work, Bianchi et al. (2023) show bias amplification in occupations and personal attributes but examine only 20 manually curated prompts. We provide a larger-scale, automated analysis offering broader, more precise insights. Cheong et al. (2024) study gender and racial bias in occupations and report stronger gender imbalance than U.S. labor statistics; we confirm and extend this by analyzing activities as well. Luccioni et al. (2024) measure bias without explicit gender/race labels, identifying distributional differences across 4,380 images from 146 occupation prompts, and explicitly call for deeper studies, which we conduct in this work. Similarly, Cho et al. (2023) observe differences in 737 images from 83 occupation prompts using gendered vs. gender-neutral phrasing. In contrast, we sample more images, filter and crop unsuitable ones, and consider scenarios beyond occupations.
Previously, Zhang et al. (2024c); Wu et al. (2024); Mannering (2023) investigate gender bias in person–object co-occurrence, but on narrow object sets (mainly clothing). We demonstrate gender bias in contexts such as traditional household roles with substantial societal relevance. Ungless et al. (2023) show poor representation of non-binary identities; we exclude them due to conceptual and technical limitations (Appendix J). Likewise, Ghosh & Caliskan (2023) show weak representation of national identities and overly sexualized depictions of women, especially for Global South prompts; Jha et al. (2024) quantify national stereotyping. Wu et al. (2024) generate 800,000 images from 200,000 gender-neutral prompts, but few images per prompt limit conclusions about which specific biases arise and their magnitude.
Although previous work establishes gender bias, scenarios beyond occupations remain underrepresented (Wan et al., 2024). Most studies are small-scale, typically under 200 prompts with no more than 20 images per prompt. Even larger benchmarks (Luo et al., 2024; 2025) still center on occupation prompts. Accordingly, we answer calls for in-depth analyses across a broader range of scenarios (Wan et al., 2024; Luccioni et al., 2024) and document the models’ default “worldview” (De Simone et al., 2023; Katirai et al., 2024). Because many template-based prompts yield images without clearly gendered people (Lyu et al., 2025), we ensure validity by filtering images without people or depicting both men and women (see Appendix K).
@@ 3.1 Prompt Collection, Processing and Clustering
We collect prompts in four categories: (1) Activities, (2) Contexts, (3) Objects, and (4) Occupations. To study gender bias in everyday activities, we rely on the curated set from (Wilson & Mihalcea, 2017), i.e. Activities. The activities in (Wilson & Mihalcea, 2017) were gathered through Amazon Mechanical Turk from U.S. based workers, who were asked to provide short phrases describing recent activities. Including these 1405 activities provides insight into how stereotypical gender roles in everyday life are reflected in T2I models. Analyzing gender associations with contexts and objects further extends this analysis by considering everyday situations beyond specific activities. To study gender bias in relation to people and places, i.e. Contexts, we collect a set of 737 scene classes from the SUN Database (Xiao et al., 2010; 2016). We include all classes but do not consider fine-grained distinctions, e.g. “inside” the church and “outside” the church leads to the context “church”.
We collect 500 common physical objects from WordNet Fellbaum (1998), i.e. Objects. To select these 500 objects, we filter all noun hyponyms of the object.n.01 WordNet synset using a list of the most common English words. Additionally, we manually remove a small number of unsuitable synsets, e.g. synsets that refer to people or body parts, or places that are already covered in the contexts prompt group. We retain the top 1000 most frequent lemmas and re-rank them based on their concreteness following Brysbaert et al. (2014). Finally, we select the top 500 most concrete lemmas from the top 1000 most frequent WordNet lemmas. We also include occupations to align with prior work and to examine occupation-related gender bias in T2I models. Our occupation list is comparatively large, as we include all 575 occupations listed by Bureau of Labour Statistics (2023) rather than a subset, i.e. Occupations.
Using Yi-1.5-34B (Young et al., 2024) (comparison in Section C.2), we convert collected activities, contexts, objects, and occupations into syntactically coherent prompts. Prompts are gender-neutral and begin with “a person”. We apply group-specific templates (Table 1); the LLM prompts used to fill them are in Section C.1. We also simplify occupation descriptions, as (Bureau of Labour Statistics, 2023) is often overly detailed. We generate 5 variations per prompt by replacing the prefix “a person” with “an individual”, “someone”, “a friend”, and “a colleague”, which do not include any gender information. Prompt variations increase diversity and are essential for valid analysis of T2I social bias (Seshadri et al., 2022; Sclar et al., 2024; Hida et al., 2024). Section D.4 shows variations do not lead to significant skew towards male- or female-gendered images.
Activity a person who is {{verb}}-ing {{activity}}” Context a person {{in/on/at…}} the {{context}}”
Objects a person and {{a object}}” Occupations a person working as {{occupation}}”
@@
For a concise presentation of our findings from the 3217 prompts, we cluster prompts in all four prompt groups into semantically coherent concepts. We apply the following variation of BERTopic (Grootendorst, 2022) to cluster prompts. First, we embed all prompts using a sentence embedding model (Reimers & Gurevych, 2019). Then, we reduce the embedding dimensions to 16 using UMAP (McInnes et al., 2018). Using HDBSCAN clustering (McInnes & Healy, 2017) with cosine distance, we determine the concept clusters, adding unclustered prompts by HDBSCAN to the cluster with the nearest centroid. The detailed settings are in Section C.3. We arrive at 165 activity clusters, 91 context clusters, 62 object clusters, and 76 occupation clusters.
After clustering prompts, we summarize all clusters by an LLM (see Section C.3). Summarizing a list of prompts requires more in-depth understanding and reasoning than prompt processing, so we use Llama-3-3-70B-Instruct based on manual inspection. A comparison to other LLMs and our exact prompt template is in Section C.3. For the purposes of our analysis, we further merge clusters with the same summary (e.g. different variants of shopping-related activities), as they would be indistinguishable for the reader. However, our code release will include the fine-grained clusters.
@@ 3.2 Image Generation and Gender Identification
We generate images using 5 models, which represent the state-of-the-art among open models at the time of writing: (1) Flux (Black Forest Labs, 2024), (2) Flux-Schnell, (3) Stable Diffusion 3.5 Large (Stability AI, 2024), (4) Stable Diffusion 3.5 Medium, and (5) Stable Diffusion 3 Medium (Esser et al., 2024). These are latent diffusion models (Rombach et al., 2022) based on Diffusion Transformers (Peebles & Xie, 2023). Other recent strong models, such as Flux-Krea and Qwen-Image (Wu et al., 2025), were concurrently released to this study and unavailable when conducting experiments. For each combination of model and prompt, we generate 40 images per prompt variation. This results in 5 models $\times$ 5 prompt variations $\times$ 3217 prompts $\times$ 40 images = 3,217,000 images.
To analyze gender bias in generated images, we identify the perceived gender of the people shown in the images using a two-step process. First, we detect all people in the images using the object detector YOLOv10 (Wang et al., 2024) and obtain bounding boxes for the detected individuals. Next, we crop each detected person’s bounding box and pass it to an MLLM, InternVL2-8B (Chen et al., 2024b; c), along with a prompt asking the model to identify the person’s gender as “female”, “male”, or “unclear/cannot tell”. We focus on binary perceived gender, specifically men and women, for several reasons. First, it is unclear whether non-binary gender has distinct visual representation, in any case current T2I models do not generate features that clearly indicate non-binary gender. Second, current MLLMs do not consider non-binary gender as an option, as shown in Section E.2. While not addressed in this paper, the fact that models output gender as binary is a separate issue that warrants further discussion.
Using bounding boxes instead of the full image helps mitigate bias from the person’s context, i.e. predicting the gender based on the background and not the person’s features, as MLLMs also exhibit gender bias (Girrbach et al., 2025). It also prevents confusion when multiple people of potentially different genders appear in the image. In Section E.1, we provide the detailed prompt used for gender identification and verify, using human-labeled data, that the MLLM used in this study can identify perceived gender with near-perfect accuracy. It is important to note that assigning a person’s gender can be problematic, because it cannot necessarily be perceived from an image, and gender is a spectrum. Therefore, good practice with images of real people is to have people self-identify their gender. However, T2I models create images that are not real, so assigning perceived gender to the images is more acceptable as there is no risk of misidentifying a real person.
@@
Table 2: Prompt groups with remaining prompts and images in brackets after filtering.
@@ 3.3 Image and Prompt Filtering
We exclude images and prompts that are not suitable for analyzing gender bias, i.e. if (a) There is no person in the image; (b) There is no person in the image whose gender can be clearly identified (see Section E.3 for details); (c) There are multiple people in the image and there is at least one man and one woman. If there are multiple people, we keep the image if people with all the same gender are shown or where the gender of other people is labeled “unclear/cannot tell.” (for example people in the background). Additionally, we exclude entire prompts for a model if fewer than 100 out of 200 images remain after filtering. Since we analyze gender bias at the prompt level, we can only consider prompts where we can reliably estimate the ratio of male and female people in the images. The number of remaining prompts for each model is in Table 2.